Y=ƒ(2x-1)/(x^2+1)ƒ(x)= sin(x^2 )Then dy/dx .
..{note x^2 means xsquare}

Grade:

1 Answers

Ramesh V
14 years ago

Taking u r question to be :

f(x) = sin(x2) and y=f(2x-1)/(x2+1)

so y = sin((2x-1)2) / (x2+1)

dy/dx ={ 4.(2x-1).(x2+1).cos((2x-1)2) _ 2x.sin((2x-1)2) } / (x2+1)

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it
here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE
preparation. All the best.

Regards,
Ramesh